Hi. I've solved for t but still can't figure out how to solve for the questions. Thank you for your help.
Here is the problem:
What is the slope of the line x=-3t+10 and y=4t+2.
(a) m=
(b) Where does the line cross the x-axis? (______,0)
(c) Where does the line cross the y-axis? (0,______)

Answer by stanbon(75887)

Here is the problem:
Comment: Didn't notice the problem was parametric.
------
What is the slope of the line x=-3t+10 and y=4t+2.
3t = -x+10
t = (-x+10)/3
-----
4t = y-2
t = (y-2)/4
----
Equate the "t" values:
(-x+10)/3 = (y-2)/4
Cross-multiply::::
3(y-2) = 4(-x+10)
3y-6 = -4x + 40
3y = -4x + 46
y = (-4/3)x + (46/3)
---------------------------
--------------------------
(a) m= -4/3
--------------------
(b) Where does the line cross the x-axis? (______,0)
y = (-4/3)x + (46/3)
(-4x+46)/3 = 0
-4x = -46
x = 46/4
-------------------
(c) Where does the line cross the y-axis? (0,______)
y = (-4/3)x + (46/3)
y = (46/3)
==============================
Cheers,
Stan H.
==============